OpenStudy (kkutie7):

Doing another Integration:

OpenStudy (kkutie7):

\[\int 3^{x}e^{x}\] I'm not sure where to start

OpenStudy (kkutie7):

this is what I was thinking first: \[\int e^{xln{3}}e^{x}dx\] then what u sub for xln3?

OpenStudy (shubhamsrg):

Have you learnt integration by parts ?

OpenStudy (kkutie7):

yes I have

ganeshie8 (ganeshie8):

this should do : \(a^m\cdot a^n = a^{m+n}\)

OpenStudy (shubhamsrg):

That'd also do. Yes.

OpenStudy (shubhamsrg):

Otherwise it's a simple application of integrating by parts.

OpenStudy (kkutie7):

\[\int e^{xln3+x}dx\]

OpenStudy (kainui):

@shubhamsrg Integration by parts here is a bad idea

OpenStudy (kkutie7):

\[\frac{1}{ln(3)+1}\int e^{u}du\]

OpenStudy (shubhamsrg):

I = int(e^x 3^x) = e^x 3^x - int(e^x 3^x ln 3) I= e^x 3^x - I ln3 It's not too difficult to calculate I from here I guess.

OpenStudy (kkutie7):

answer=\[\frac{3^{x}e^{x}}{ln{3}+1}\] right?

OpenStudy (anonymous):

\[\int\limits 3^x e^x dx=3^x.e^x- \int\limits3^x \ln(3) e^xdx+C\]\[(1+\ln(3))\int\limits 3^x e^x dx=3^x.e^x+C\]\[\int\limits 3^x e^x dx=\frac{3^xe^x}{1+\ln(3)}+C'\] Looks fine to me using by parts

ganeshie8 (ganeshie8):

that looks good ! below is, i think, a slightly better algebra.. \[\int 3^xe^x\,dx = \int (3e)^x\, dx = \int e^{x\ln (3e)}\, dx = \dfrac{e^{x\ln(3e)}}{\ln(3e)}+C = \dfrac{(3e)^x}{\ln(3e)}+C\]

OpenStudy (kkutie7):

Yes I could have simplified it more. I also forgot to add the constant.

OpenStudy (kainui):

\[\int b^x dx = \frac{b^x}{\ln b}\] \[\int 3^xe^x dx = \int e^{(1+\ln 3)x}dx \] \[b=e^{1+\ln 3}\]